a bar magnet of length 3cm has points A and B along its axis at distances of 24cm and 48cm on the opposite sides. how to calculate the ratio of magnetic field at these points ?

Grade:11

2 Answers

Anu
6 years ago
B is inversely proportional to one upon (distance)3 this you see in on axial and normal bisector (magnetic field due to bar magnet)Now, Ba/Bb=(48/24)3 =8/1 answer
ankit singh
3 years ago
Length of magnet = 3 cm
Distance between point a and center of magnet= 24 cm
Distance between point b and center of magnet= 48 cm
The magnetic field is
B=\dfrac{\mu_{0}\times 2m}{4\pi\times d^3}
B\propto\dfrac{1}{d^3}
The magnetic field is inversely proportional to the cube of the distance.
We need to calculate the ratio of magnetic fields at these points
Using formula of magnetic field
\dfrac{B_{a}}{B_{b}}=\dfrac{d_{b}^3}{d_{a}^2}
\dfrac{B_{a}}{B_{b}}=\dfrac{48^3}{24^3}
\dfrac{B_{a}}{B_{b}}=\dfrac{8}{1}